OpenStudy (anonymous):

limit as x approaches zero of quantity negative six plus x divided by x to the fourth power. Would the limit be 0 or does not exist?

OpenStudy (zale101):

\[\lim_{x \rightarrow 0}~[-6+\frac{x}{x^4}]\] Like this?

OpenStudy (anonymous):

No @Zale101

OpenStudy (zale101):

How's it like then?

OpenStudy (anonymous):

\[\frac{ -6+x }{ x^4 }\]

OpenStudy (anonymous):

^ @Zale101

OpenStudy (zale101):

Oh i see \[\lim_{x \rightarrow 0} ~[\large \frac{-6+x}{x^4}]= \\lim_{x \rightarrow 0} ~[\large \frac{\frac{-6}{x^4}+\frac{x}{x^4}}{\frac{x^4}{x^4}}]= \lim_{x \rightarrow 0} ~[\large \frac{\frac{-6}{x^4}+\frac{1}{x^4}}{{1}}]=\]

OpenStudy (zale101):

Now, what happens if i sub in x=0?

OpenStudy (anonymous):

It's 0/1 which is undefined? @Zale101

OpenStudy (zale101):

\[\lim_{x \rightarrow 0} ~[\large \frac{\frac{-6}{x^4}+\frac{1}{x^3}}{1}]=\large \frac{\frac{-6}{0}+\frac{1}{0}}{1}\] Therefore, it does not exist because something over a zero is indeterminate.

OpenStudy (anonymous):

Ok thank you!

OpenStudy (anonymous):

@Zale101 Could you help me with a couple more?

OpenStudy (zale101):

Sure.